Probability questions
692 published questions with worked solutions.
- If X B(35,p) such that 7P(X=0)=P(X=1) then the value of P(X=15) P(X=20) is equal to
- Let X and Y are two events such that P(X Y) = P(X Y). Statement 1: P(X Y') = P(X' Y) = 0 Statement 2: P(X) + P(Y) = 2P(X Y)
- If six students, including two particular students A and B, stand in a row, then the probability that A and B are separated with...
- There are two balls in an urn. Each ball can be either white or black. If a white ball is put into the urn and there after a ball...
- Three numbers are chosen at random without replacement from 1, 2, 3, . . . . 8 . The probability that their minimum is 3 , given...
- The probability of a man hitting a target is 25. He fires at the target k times (k, a given number). Then the minimum k, so that...
- Given two independent events, if the probability that exactly one of them occurs is 26 49 and the probability that none of them...
- A, B, C try to hit a target simultaneously but independently. Their respective probabilities of hitting the targets are 34, 12,...
- If the events A and B are mutually exclusive events such that P(A) = 3x + 1 3 and P(B) = 1 - x 4 , then the set of possible...
- Let A and B be two events such that P ( A B ) = 16, P(A B) = 14 and P ( A ) = 14, where A stands for the complement of the event...
- If A and B are two events such that P(A B) = P(A B), then the incorrect statement amongst the following statements is :
- A set S contains 7 elements. A non-empty subset A of S and an element x of S are chosen at random. Then the probability that x A...
- Let A and E be any two events with positive probabilities. Statement I: P(E/A) P(A/E)P(E). Statement II: P(A/E) P(A E).
- Suppose that Box I contains 6 red balls and 9 green balls, and Box II contains 8 red balls and 12 green balls. All the balls of...
- Two cards are drawn at random from a pack of 52 cards. The probability of getting at least a spade and an ace is
- A five-digit number is written down at random. The probability that the number is divisible by 5 and no two consecutive digits...
- If the letters of the word ATTEMPT are written down at random, the chance that all Ts are consecutive is
- In a single cast with two dice the odds against drawing 7 is
- Let x = 33^ n . The index n is given a positive integral value at random. The probability that the value of x will have 3 in the...
- Three dice are thrown simultaneously. The probability of getting a sum of 15 is
Other chapters
Alcohols, Phenols and EthersAldehydes, Ketones and Carboxylic AcidsAlternating CurrentAminesAnatomy of Flowering PlantsAncient Indian HistoryAnimal KingdomApplication of DerivativesArea Under CurveAromatic Compounds and Aryl HalidesAtomsBasic Mathematics and LogarithmBinomial TheoremBiodiversity and ConservationBiological ClassificationBiomoleculesBiotechnology and Its ApplicationsBiotechnology: Principles and ProcessesBody Fluids and CirculationBreathing and Exchange of GasesCell Cycle and Cell DivisionCell: The Unit of LifeChemical Bonding and Molecular StructureChemical Coordination and Integration